OpenStudy (anonymous):

use f(x)=x^2 and g(x)=x+5. Find the function and simplify the function rule. (f o g)(x)

OpenStudy (anonymous):

@Nnesha @triciaal

OpenStudy (triciaal):

f o g(x) use the value for g(x) as x in f(x)

OpenStudy (anonymous):

so f(g(x)) ----> f(x+5). do i multiply f(x) to the g(x)?

OpenStudy (xapproachesinfinity):

think about it as this \(f(x)=x^2\) i can replace that x with something else like \(f(\square)=(\square)^2\) so \(fog(x)=f(g(x))=(g(x))^2\)

OpenStudy (triciaal):

f(x)=x^2 and g(x)=x+5. f(x) = x^2 and x = (x + 5) so it becomes (x + 5)^2

OpenStudy (anonymous):

so i get x^2+25?

OpenStudy (triciaal):

incomplete try again multiply each term

OpenStudy (xapproachesinfinity):

Not quite! recall that \((a+b)^2=a^2+2b+b^2\)

OpenStudy (xapproachesinfinity):

forgot a in there

OpenStudy (triciaal):